OpenStudy (anonymous):

What is its maximum height? Its initial vertical speed is 8.8 m/s and the acceleration of gravity is 9.8 m/s 2 . Neglect air resistance. Answer in units of m

OpenStudy (anonymous):

Maximum height occurs at v = 0 m/s. Vf^2 = Vi^2 + 2a(Xf - Xi) 0 = 8.8^2 + (2)(-9.8)(Xf - 0) 0 = 8.8^2 - 19.6Xf 19.6 Xf = 8.8^2 Xf = 3.95 m

OpenStudy (anonymous):

Thank you so much!!